# RNBF: Real-Time RGB-D Based Neural Barrier Functions for Safe Robotic Navigation

### v1

@@ IV-A5 Continuity & Differentiability of Neural SDF
Because $h_{\text{sdf}}(\xi)$ is composed of continuous functions (positional embedding, affine transformations, $\operatorname{Softplus}$), it is continuous on $\mathbb{R}^{3}$. Moreover, each component is differentiable, so by the chain rule, the gradient $\nabla h_{\text{sdf}}(\xi)$ exists and remains continuous everywhere [28, 29, 30]. Concretely,
@@ IV-B RNBF Pipeline
Unlike existing work that twists standard CBF-QP formulations (Eq. (8)) to accommodate noises and instability in the learned SDFs, we are able to combine SDF outputs from the vision pipeline smoothly with existing reactive safe controller, like CBF-QP, since our SDF learning process has explicitly considered sensor noises. Assuming robot state $x$ contains robot position, i.e. $x=[\xi;\xi^{\prime}]$ and $\xi^{\prime}$ are robot state elements that excludes 3D position, the idealized CBF-QP in (8) can be rewritten as:
$u_{\text{cbf}}=\operatorname*{arg\,min}_{{u}\in\mathbb{R}^{m}}\frac{1}{2}||u-u_{\text{nom}}||_{2}^{2}$ (19)
@@
As stated in Section II, we reasonably assume that $h_{\text{sdf}}(\xi_{\text{rob}})=h_{\text{sdf}}(\xi_{\text{cam}})=h_{\text{sdf}}(\xi)$. It is noteworthy that while given any control affine robot dynamics, $\dot{x}$ is guaranteed to be actuated, i.e. $\dot{x}$ explicitly depends on control input $u$. However, such guarantee does not necessarily hold for $\dot{\xi}$, in which case an appropriate robot model must be choosen. Our resulted outputs from the vision pipeline can be adopted by any robot safe navigation algorithms that plan collision-free trajectories given real-time distance and the surface normal (or gradient) to the nearest obstacle.
@@ V-A Experiment Setup
Our proposed approach is evaluated on a Fetch robot running at 20 Hz in simulation and robot hardware. For simulations, we employ a realistic hospital environment in Gazebo, as shown in Figure 4, and define four scenarios with different initial and target combinations. For real-world validation, we equip the Fetch robot with an Intel RealSense D435i RGB-D camera (Figure  1) and utilize a motion capture system to acquire real-time camera poses. All experiments are run on an Intel Core i7-12700K CPU with a GeForce RTX 3090 Ti GPU.
Figure 4: Overhead view of the environment with planned trajectories.
@@ V-B Robot Dynamics
Nominal Controller: Our nominal controllers are designed to approximatedly follow linear trajectories $u_{\text{nom}}^{l}$ pointing from the robot’s position $\xi$ to the target $\xi^{*}$ with a speed of 0.5 $m/s$, where $\psi$ is the angle difference between the current robot pose $\theta$ and the desired robot pose estimated using the orientation of $u_{\text{nom}}^{l}$, and $\Delta t$ is the update timestep of the nominal controller. The approximation is not exact due to the non-holonomic property of differential drive systems.
$u_{\text{nom}}^{l}=0.5\frac{\xi-\xi^{*}}{||\xi-\xi^{*}||_{2}}\quad v_{\text{nom}}=||u_{\text{nom}}^{l}||_{2}\quad\omega_{\text{nom}}=\frac{\psi}{\Delta t}$ (24)
@@ V-C Effect of Depth Noise on SDF
When deploying our pipeline on a real robot, we use a low-cost Intel RealSense D435i sensor, whose depth estimates are notably noisy. Such noises introduce spurious artifacts that would degrade the SDFs and their gradients from the neural network if left attended. Such artifacts will disqualify the SDFs from being approximate barrer functions. To mitigate this issue, we adopt a huber loss for $L_{\text{near\_surf}}$ (12), which reduces the impact of outliers while still enforcing tight supervision near the surface. Figure 6 compares SDF reconstructions using RealSense depth data. The L1 loss based model amplifies sensor noise, leading to rough surfaces and unstable gradients, whereas the Huber-based model produces smoother geometry and more stable gradients. The choice of $\delta$ significantly influences reconstruction quality: smaller values of $\delta$ suppress noise more effectively but risk oversmoothing fine details, while larger values retain geometric features but may allow noise to propagate. In our real-world experiments, we find $\delta=[0.01,0.05]$ to offer a favorable balance, though it remains a tunable parameter depending on the quality of depth sensor. Although we do not claim perfect fidelity, the result is sufficiently accurate for our pipeline. Overall, the huber loss preserves the key property of a smoothly varying SDF, essential for real-time CBF-based control under imperfect depth measurements.
@@ V-D Experiment Results
Figure 7: Trajectory Comparison: RNBF-CBF-QP vs. Idealized CBF-QP in Scenario 1.
In this section, we present the experimental results that highlight the effectiveness of our proposed RNBF-Control pipeline. RNBF-CBF-QP is compared against an idealized baseline that assumes full knowledge of all obstacle locations and geometries. We display in Figure 7 the trajectories generated by RNBF-CBF-QP, the idealized CBF-QP and the nominal controller, and in Figure 8 multiple RNBF-CBF-QP runs overlaid for scenario 1.
In Table I, the performances of running Scenario 1 for 10 times using RNBF-CBF-QP and idealized CBF-QP are analyzed quantitatively using the average time (in seconds) that the robot takes to travel from the initial to the target locations (durations), the percentage of collision-free trajectories among the repeated ones (safe %), the chance for the robot to reach the target (reached %). Both RNBF-CBF-QP and the idealized CBF-QP achieved a 100% collision-free rate and a 100% success rate in reaching the destination. RNBF-CBF-QP takes a little more time (36.7 seconds) than the idealized CBF-QP (29.0 seconds) because it relies solely on RGB-D sensor data for constructing barrier functions, rather than using prior knowledge of obstacle placements.
Beyond gazebo simulations, the proposed pipeline is further validated in real-world experiments involving both static and quasi-static obstacles (Figure 1). RNBF-CBF-QP demonstrates reactive obstacle avoidance ability when facing both obstacle types under partial observability and sensor noise. We provide supplementary videos showing these trials, which reveal that our method is capable of accurately building and updating the barrier function on-the-fly, thereby ensuring safety in novel environments.
@@
Idealized CBF-QP 100 100 29.0
TABLE I: Performance comparison of RNBF-CBF-QP with Idealized CBF-QP in Scenario 1.
Figure 8: Overlaid RNBF-CBF-QP Trajectories from multiple runs in Scenario 1.
@@ VI CONCLUSIONS
The paper introduces a novel pipeline for real-time CBF construction in unknown environments using affordable sensors like RGB-D cameras. Our framework is robust to sensor noise while generating a smooth and differentiable neural SDF representation, thereby preserving the core assumptions required by standard CBFs. Experiments in both simulation and on a physical Fetch robot confirm that our approach enables safe, reactive navigation without relying on expensive LiDAR sensors and can readily integrate with established SDF-based controllers. Moreover, the system’s modular design allows it to serve as a dedicated safety layer in broader robotic applications, given only onboard RGB-D sensing.

### v3

@@ IV-A4 Continuity & Differentiability of Neural SDF
Because $h_{\text{sdf}}(\xi)$ is composed of continuous functions (positional embedding, affine transformations, $\operatorname{Softplus}$), it is continuous on $\mathbb{R}^{3}$. Moreover, each component is differentiable, so by the chain rule, the gradient $\nabla h_{\text{sdf}}(\xi)$ exists and remains continuous everywhere [hornik1989multilayer, park2019deepsdf, tancik2020fourier]. Concretely,
@@ IV-B RNBF Pipeline
Unlike existing work that twists standard CBF-QP formulations (Eq. (7), (8)) to accommodate noises and instability in the learned SDFs, we are able to combine SDF outputs from the vision pipeline smoothly with existing reactive safe controller, like CBF-QP, since our SDF learning process has explicitly considered sensor noises. Assuming robot state $x$ contains robot position, i.e. $x=[\xi;\xi^{\prime}]$ and $\xi^{\prime}$ are robot state elements that exclude 3D position, the CBF-QP in (7), (8) can be rewritten as:
[…]
@@
As stated in section II, we reasonably assume that $h_{\text{sdf}}(\xi_{\text{rob}})=h_{\text{sdf}}(\xi_{\text{cam}})=h_{\text{sdf}}(\xi)$. It is noteworthy that while given any control affine robot dynamics, $\dot{x}$ is guaranteed to be actuated, i.e. $\dot{x}$ explicitly depends on control input $u$. However, such guarantee does not necessarily hold for $\dot{\xi}$, in which case an appropriate robot model must be chosen. Our resulted outputs from the vision pipeline can be adopted by any robot safe navigation algorithms that plan collision-free trajectories given real-time distance and the surface normal (or gradient) to the nearest obstacle.
@@ V-A Experiment Setup
Our proposed approach is evaluated on a Fetch robot running at 20 Hz in simulation and robot hardware. For simulations, we employ a realistic hospital environment in Gazebo, as shown in 4(a), and define four scenarios with different initial and target combinations. For real-world validation, we equip the Fetch robot with an Intel RealSense D435i RGB-D camera (Figure 1) and utilize a motion capture system to acquire real-time camera poses. All experiments are run on an Intel Core i7-12700K CPU with a GeForce RTX 3090 Ti GPU.
Figure 4: (a) Simulation environment setup with four different scenarios. (b) Ground-truth SDF used by Parametric CBF-QP. Note, RNBF-CBF-QP builds its SDF online and does not use the ground-truth SDF.
@@ V-B Robot Dynamics
Nominal Controller: Our nominal controllers are designed to approximately follow linear trajectories $u_{\text{nom}}^{l}$ pointing from the robot’s position $\xi$ to the target $\xi^{*}$ with a speed of 0.5 $m/s$, where $\psi$ is the angle difference between the current robot pose $\theta$ and the desired robot pose estimated using the orientation of $u_{\text{nom}}^{l}$, and $\Delta t$ is the update timestep of the nominal controller. The approximation is not exact due to the non-holonomic property of differential drive systems.
[…]
@@ V-C Effect of Depth Noise on SDF
When deploying our pipeline on a real robot, we use a low-cost Intel RealSense D435i sensor, whose depth estimates are notably noisy. Such noises introduce spurious artifacts that would degrade the SDFs and their gradients from the neural network if left attended. Such artifacts will disqualify the SDFs from being approximate barrier functions. To mitigate this issue, we adopt a huber loss for $L_{\text{near\_surf}}$ (11), which reduces the impact of outliers while still enforcing tight supervision near the surface. Figure 6 compares SDF reconstructions using RealSense depth data. In this work, sensor noise refers specifically to the depth noise inherent in affordable RGB-D sensors. The L1 loss based model amplifies sensor noise, leading to rough surfaces and unstable gradients, whereas the huber-based model produces smoother geometry and more stable gradients. The choice of $\delta$ significantly influences reconstruction quality: smaller values of $\delta$ suppress noise more effectively but risk oversmoothing fine details, while larger values retain geometric features but may allow noise to propagate. In our real-world experiments, we find $\delta=[0.01,0.05]$ to offer a favorable balance, though it remains a tunable parameter depending on the quality of depth sensor. Although we do not claim perfect fidelity, the result is sufficiently accurate for our pipeline. Overall, the huber loss preserves the key property of a smoothly varying SDF, essential for real-time CBF-based control under imperfect depth measurements.
@@ V-D Experiment Results
Figure 7: Scenario 1 (simulation): Trajectory comparison. Dashed curves represent trajectories, with mean values shown where applicable. For RNBF-CBF-QP, a shaded tube depicts the variance of the mean trajectory over 10 rollouts. Parametric CBF-QP, which assumes perfect obstacle knowledge, produces nearly identical paths across runs; only one representative trajectory is plotted for clarity. iSDF-CBF-QP is likewise shown with a single representative run since it collided in all trials. The nominal controller is deterministic and follows the shortest path from start to goal.
In this section, we present the experimental results that highlight the effectiveness of our RNBF-based safety filter. While in this work we demonstrate its use with a CBF-QP controller (RNBF-CBF-QP), the RNBF-based safety filter pipeline can be readily adapted to any SDF-based controller or safety filter that consumes an SDF and its gradients. RNBF-CBF-QP is compared against two baselines: (i) Parametric CBF-QP, which assumes full knowledge of all obstacle locations and geometries using the ground-truth SDF of the environment (4(b)), and (ii) iSDF-CBF-QP, which integrates the original iSDF reconstruction into a CBF-QP without floor segmentation or depth-denoising. We display in Figure 7 the trajectories generated by RNBF-CBF-QP, the parametric CBF-QP, iSDF-CBF-QP and the nominal controller for scenario 1.
@@
Parametric CBF-QP 100 100 29.0
iSDF-CBF-QP 0 0 -
TABLE I: Performance comparison of RNBF-CBF-QP with Parametric CBF-QP & iSDF-CBF-QP in Scenario 1 over 10 independent trials.
Table I reports quantitative results for ten independent runs of Scenario 1 with RNBF-CBF-QP, Parametric CBF-QP, and iSDF-CBF-QP. Performance is measured by the average traversal time from start to goal (duration, s), the proportion of collision-free trajectories (safe %), and the proportion of successful goal reaches (reached %). Both RNBF-CBF-QP and the Parametric CBF-QP achieved a 100% collision-free rate and a 100% success rate in reaching the destination. However, RNBF-CBF-QP requires slightly more time to reach the goal (36.7 s vs. 29.0 s) owing to its exclusive reliance on online RGB-D sensing to construct barrier functions, as opposed to the Parametric CBF-QP’s use of full prior knowledge of obstacle locations. In contrast, the iSDF-CBF-QP fails in all trials, with $0\%$ success rate in reaching the destination and repeated collisions with the obstacles. Two factors primarily drive this failure: (1) Floor misclassification: without a segmentation module, large planar floor regions are incorrectly treated as obstacles, preventing the controller from finding a feasible path; this issue arises in both simulation and real-world deployments. (2) Depth-noise artifacts: sensor noise introduces spurious structures into the neural SDF (section V-C), yielding unstable gradients and erratic control commands; while less relevant in simulation due to noise-free gazebo depth sensing, it significantly impacts physical robot navigation (Figure 9). RNBF-CBF-QP mitigates both issues through explicit floor segmentation and huber-loss based near-surface supervision, enabling stable gradients and reliable, collision-free navigation.
Figure 8: Real-world experiments. Shown are collision-free RNBF-CBF-QP trajectories overlaid on video frames for static (single obstacle - shape 1 & 2, and multi obstacle) and quasi-static obstacle scenarios.
Beyond gazebo simulations, the proposed pipeline is further validated in real-world experiments involving both static and quasi-static obstacles (Figure 1 & Figure 8). For both types of obstacles, RNBF-CBF-QP reached the goal collision-free under partial observability and depth noise. In contrast, iSDF-CBF-QP exhibited the same failure modes as in simulation leading to collisions and failure to reach the goal as shown in Figure 9. We provide supplementary videos [^1] Videos and supplementary material are available at https://satyajeetburla.github.io/rnbf/, which reveal that our method is capable of accurately building and updating the barrier function on-the-fly, thereby ensuring safety in novel environments.
Figure 9: Trajectory comparison of RNBF-CBF-QP vs iSDF-CBF-QP for real-world deployment with static obstacle.
@@ VI CONCLUSIONS
The paper introduces a novel pipeline for real-time CBF construction in unknown environments using affordable sensors like RGB-D cameras. Our framework is robust to sensor noise while generating a smooth and differentiable neural SDF representation, thereby preserving the core assumptions required by standard CBFs. Experiments in both simulation and on a physical Fetch robot confirm that our approach enables safe, reactive navigation and can readily integrate with established SDF-based controllers. Moreover, the system’s modular design allows it to serve as a dedicated safety layer in broader robotic applications, given only onboard RGB-D sensing.